OpenStudy (russianboss1):

which answer is equal to ^6sqrt(729)^6 a: -3 b:-729 c:729 d:3 I know it is probably either b or c

OpenStudy (darrien117):

i would have to say c

OpenStudy (astrophysics):

\[\huge \sqrt[6]{729^6}\] \[\huge \sqrt[n]{x^m} = x^{\frac{ m }{ n }}\]